### v1

# Motional entanglement in low-energy collisions near shape resonances

Abstract
Einstein, Podolsky, and Rosen discussed their paradox in terms of measuring the positions or momenta of two particles. These can become entangled upon scattering, but how much entanglement can be created in this process? Here we address this question with fully coherent calculations of bipartite scattering in three-dimensional space, quantifying entanglement by the inverse of the single particle purity. We show that the standard plane-wave description of scattering fails to capture the entanglement properties, due to the essential role of quantum uncertainty in the initial state. For a more realistic description of a scattering setup and narrow initial momentum dispersion, we find the entanglement to scale linearly with the scattering cross section, including strong enhancement close to shape resonances. We discuss how the generation of motional entanglement can be detected in experiment. Our results open the way to probing and eventually using entanglement in quantum collisions.
Introduction
Over the past few decades, entanglement has become the cornerstone of quantum information science [1]. The puzzling nature of entanglement as “spooky action at a distance” is exemplarily captured by the Einstein-Podolsky-Rosen (EPR) paradox [2]. In the original EPR gedankenexperiment, it is the positions, or momenta, of two particles that are entangled. Such entanglement could be leveraged for quantum technologies such as quantum sensing [3, 4], but also for the coherent control of reactive collisions [5, 6]. Entanglement requires an interaction between the relevant degrees of freedom which, for motional entanglement, can be leveraged in breaking up a composite object as happens in dissociation [7, 8] or photoionization [9], or it can be picked up in collisions. While collisions were shown to be a source of entanglement for internal degrees of freedom [10, 11, 12, 13], their use for generating motional entanglement remains largely unexplored.
This may in part be due to the difficulty to characterize entanglement in the infinite-dimensional Hilbert spaces of continuous variables (CV) [14]. Identification of CV entanglement in general is limited to bi-partite pure states, using quantifiers such as the von Neumann entropy [15] or the single-particle purity [16] which are based on the reduced density matrices of each particle. Alternative criteria such as the positive partial transpose (PPT) [17, 18] or quantifiers based on the Shannon entropy [19] or the Husimi $Q$ distribution [20] are necessary and sufficient only for Gaussian states, as encountered for photons [21], but are unsuitable to quantify entanglement generated in collisions of free particles.
A second difficulty arises from the challenge of a full quantum-mechanical description of free particle collisions in three-dimensional space [22]. To date, the generation of entanglement in collisions was studied only in the limit of pure $s$-wave scattering [23] and in one spatial dimension [8, 24, 25, 26, 27, 28]. Confinement to one spatial dimension leads to orthogonality of the transmitted and reflected amplitudes, making scattering in one and more spatial dimensions fundamentally different. Analysis in 3D [23] suggested a connection of the entanglement generated in a collision and the collision cross section, at least for $s$-wave collisions and under conditions where the phase shift as a function of the collision energy remains nearly constant. However, pure $s$-wave scattering misses key features of quantum collisions — partial wave interference and scattering resonances. The latter lead to peaks in the cross section. Will this carry over to entanglement? In other words, can scattering resonances be used to maximize the generation of entanglement in collisions?
Here, we answer this question affirmatively. To this end, we consider elastic collisions between two distinguishable spinless particles in three-dimensional space and quantify the entanglement generated upon the collision using the inverse single-particle purity [24]. We show that, in the standard plane-wave description, no entanglement is generated. Solving the scattering problem for initial states with finite uncertainty, we examine the temporal evolution of entanglement and analyze scattering near a variety of shape resonances and for different geometries of the incident wave packets. Our results establish a connection between motional entanglement and the collision cross section near shape resonances, provided the initial energy dispersion is sufficiently narrow. We discuss scenarios for probing the generation of entanglement in collision experiments.
Initial states with finite uncertainty
To quantify the entanglement generated upon a collision, we envision a typical collision experiment using atomic or molecular beams, where initially the particles are spatially separated. Standard quantum-mechanical scattering theory aimed at calculating cross sections [22] is only concerned with the relative motion, disregarding the center-of-mass (CM) degrees of freedom. However, when the goal is to calculate entanglement, the single-particle density matrices need to be constructed. Then the CM motion can no longer be ignored, even if relative and CM motion are not coupled. One may be tempted to assume a plane wave $e^{i\vec{P}_{0}\cdot\vec{R}}$ for the CM motion, as was done for initially bound particles [25, 29, 30]. However, a plane wave has no momentum uncertainty, $\Delta{P}=\sqrt{\langle P^{2}\rangle-\langle\vec{P}\rangle^{2}}=0$, and an initially separable state with $\Delta{P}=0$ (or $\Delta p=0$) will remain separable after scattering, as we show in the End Matter C for any non-Coulomb potential $V(r)$. An initial state with finite uncertainty is also in line with our physical intuition about collision experiments which start with spatially separated, localized particles.
Figure 1: (a) Sketch of a collision of two structureless free particles, initially with momenta $\pm{p}_{0}\hat{z}$, positions $\mp\frac{{q}_{0}}{2}\hat{z}$ and same spatial dispersion $\vec{a}$ that spreads under time of flight. (b) Entanglement measure $\mathcal{K}$ for three initial Gaussian wave packets, and linear fit of $\mathcal{K}$ to the overlap of scattered and non-scattered wavefunctions versus time of flight over single particle mass $q_{0}=0$, $p_{0}=0.8$, under potential $V_{0}=8$ in Eq. (1)], see End Matter A for the fit parameters.
To model such an experiment, we consider head-on collisions of two particles with uncertainties in their kinetic energies and collimation, cf. Fig. 1(a). This introduces a finite dispersion also in the initial relative momenta, $\Delta p^{2}=(1/a_{x}+1/a_{y}+1/a_{z})/4$. For simplicity, we assume identical mass $m$ and opposite momenta and take the wave packets as Gaussian, so that the initial uncertainty $\Delta p\cdot\Delta r$ is minimum. The simplest potential that allows for scattering resonances is a cylindrical well with tunable width $w$ and depth $V_{0}$. The Hamiltonian is then given by
$H=\frac{\vec{P}^{2}}{4m}+\frac{\vec{p}^{2}}{m}+V(r),\quad V(r)=\begin{cases}-V_{0}/m,&\text{}r<w\\ 0,&\text{}r>w\end{cases}\,.$ (1)
Unless noted otherwise, we use $w=1$.
Entanglement generation
We use the inverse of the single-particle purity, $\mathcal{K}=1/\mathcal{P}$, where $\mathcal{P}=Tr(\rho_{1}^{2})=Tr(\rho_{2}^{2})$, to quantify the bipartite entanglement, cf. End Matter for details. When $\mathcal{K}=1$, no entanglement is generated, otherwise $\mathcal{K}>1$ (or $\mathcal{P}<1$). The finite dispersion implies a coherent superposition of different momentum components, which introduces an explicit time dependence with $\Delta r\to+\infty$ as $t\to+\infty$. The time-evolution also influences the entanglement generation as shown in Fig. 1(b), where we have used $q_{0}=0$ for clarity  [^1] For separations $q_{0}>0$, it takes time for the wave-packets to start to overlap and interact, during which the behavior of $\mathcal{K}$ is not straightforward to interpret.. Once the particles interact, $\mathcal{K}$ as a function of $t/m$ exhibits an inverse linear correlation with the overlap of the scattered ($\tilde{\Psi}$) and non-scattered ($\Psi$}) wave packets, $F(t)=|\langle{\Psi}(t)|\tilde{\Psi}(t)\rangle|^{2}=|\langle{\Psi}(0)|e^{i(H-H_{0})t}\tilde{\Psi}(0)\rangle|^{2}$, indicating that entanglement is generated by the interaction $V$. Wave packets with larger spatial dispersion $\vec{a}$ interact longer and therefore need more time for $\mathcal{K}$ to saturate. The time evolution of entanglement has largely been ignored before [25, 23, 27, 26, 28]. To regard elastic scattering as a mere redirection of incident momenta, however, violates normalization of the wave packet and energy conservation and therefore yields incorrect asymptotic values of entanglement, as we show in the End Matter B.
Figure 2: (a) Entanglement $\mathcal{K}$ vs mean incident momentum $p_{0}$, near a $p$-wave shape resonance ($p_{res}=0.87$ for $V_{0}=8$) for different spatial (momentum) dispersions $\vec{a}$ ($\Delta{p}$) and initial locations $q_{0}$. (b-c): Cross section $\sigma$ (black solid lines) and comparison to linear fits, cf. Eq. (2) [colored symbols with labels $(\Delta p,q_{0})$] vs (b) collision momentum for $V_{0}=8$ and (c) potential depth with fixed $p_{0}=0.8$.
Our subsequent analysis focuses on $\mathcal{K}(t\to+\infty)$ which characterizes the motional entanglement after the collision, shown in Fig. 2(a) as a function of the mean incident momentum $p_{0}$ for different initial conditions. For a given spatical-momentum dispersion $\vec{a}$-$\Delta{p}$ (color-coded), a larger initial separation $q_{0}$ reduces $\mathcal{K}$ (cf. different line-styles) since the wave packets meet after a long free flight and therefore their transversal overlap is small in the interaction region. This reduced entanglement generation is even more pronounced for larger initial $\Delta{p}_{x,y}$. The dependence of $\mathcal{K}$ on the initial momentum dispersion is readily understood. In the plane wave limit, $\Delta{p}\rightarrow 0$, no entanglement is generated ($\mathcal{K}\to 1$), cf. the purple curve for $\Delta{p}=0.19$ barely exceeding one. For small and intermediate $\Delta p$, the dependence of $\mathcal{K}$ on $p_{0}$ is reminiscent of the collision cross section, particularly near the $p$-wave shape resonance ($p_{res}=0.87$). For very broad momentum dispersion (cyan curve $\Delta{p}=0.79$), the resonance feature is completely smeared out by non-resonant momentum components.
Entanglement vs cross section
The close resemblance of the entanglement $\mathcal{K}$ and the collision cross section $\sigma$ as functions of collision energy for sufficiently small momentum dispersion in Fig. 2(a) suggests a linear relationship,
$\sigma_{fit}(\mathcal{K})=\gamma(\mathcal{K}-\beta)\,.$ (2)
The agreement of $\sigma_{fit}$ with $\sigma$ is excellent, except for very low energies, cf. Fig. 2(b). The fit parameters $\gamma,\beta$ are reported in Table 2. The deviation from linear behavior at low energy can be roughly explained by considering the incident momenta within the Gaussian distribution, for which $99.7\%$ of incident momenta are within $p_{0}\pm 3\Delta p$. At low collision energy, $p_{0}<3\Delta p$, such that there is a discernible amount of incident momenta that propagate in the backward direction and thus do not contribute to the collision. As $p_{0}$ increases, more momentum components contribute to the collision. Thus more entanglement is generated, explaining the monotonic increase of $\sigma_{fit}$ for small $p_{0}$ in Fig. 2(b), in contrast to the cross section which is constant. The same argument applies also to scattering with very broad momentum dispersion, for which $\mathcal{K}$ grows with $p_{0}$ monotonically without any resonance structure, cf. Fig 2(a).
To highlight that, due to resonances, stronger interaction does not necessarily imply more entanglement, Fig. 2(c) displays the close relationship between collision cross section $\sigma$ and entanglement measure $\mathcal{K}$ as a function of the potential depth $V_{0}$ (for $p_{0}=0.8$). When $V_{0}\sim 2.47$, the potential becomes deep enough to support its first $s$-wave bound state; when $V_{0}\sim 8$, a $p$-wave shape resonance forms at $p=0.8$, resulting in strong enhancement of both $\sigma$ and $\mathcal{K}$. However, for even stronger interaction, $V_{0}>8$, $\mathcal{K}$ does not increase with $V_{0}$ but decays towards $1$ (non-interacting limit). This implies that entanglement, just as the cross-section, is determined by the resonance and bound-state structure, not the strength of interaction. The strongest entanglement for wave packets with momentum dispersion $\Delta p=0.25$ is generated at $V_{0}=8.5$, where also the cross section $\sigma$ is maximized. For broader momentum dispersion, e.g. $\Delta p=0.37$, $\mathcal{K}_{max}$ shifts towards $V_{0}=8.2$ where the Wigner time-delay $\tau=d\delta_{(l=1)}/dE$ corresponding to the dwell time of the wave packet is maximal. Since the Wigner time delay is a more robust indicator of a shape resonance than the peak of the cross section in the presence of multiple partial waves and interference of different momentum components, the deviation of $\sigma_{fit}$ from $\sigma$ for $\Delta p=0.37$ in Fig. 2(c) can be attributed to more interference near the resonance than for smaller $\Delta p$.
Figure 3: Cross section $\sigma$ (black solid lines) and comparison to linear fits, cf. Eq. (2) (colored symbols) near a $p$-wave shape resonance for (a) a single potential $V_{0}=8.8$ ($p_{res}=0.64$) and (b) different potentials (Table 1) with $\sigma_{fit}$ evaluated at $p_{0}=p_{res}$ (vertical dashed lines), for $q_{0}=5$, $\vec{a}=(a_{\perp},a_{\perp},20)$, $a_{\perp}=4,10,20$, cf. Fig. 2 for the correspondence with $\Delta p$.
Table 1: $p$-wave shape resonance parameters: position $p_{res}$, resonance width (estimated by the FWHM of $\tau=d\delta_{(l=1)}/dE$), cross section $\sigma_{max}$ at $p=p_{res}$
Depth $V_{0}$ $p_{res}$ resonance width $\sigma_{max}$
8.8 0.64 0.24 100.08
8.6 0.70 0.29 83.29
8.4 0.76 0.33 70.89
8.2 0.82 0.38 61.38
8.0 0.87 0.43 53.81
7.8 0.93 0.47 47.66
7.6 0.98 0.52 42.51
7.2 1.08 0.61 34.36
The role of the resonance width is studied in Fig. 3 with Fig. 3(a) displaying the same comparison of $\sigma$ and $\sigma_{fit}$ as Fig. 2(b) but for $V_{0}=8.8$ which has a narrower resonance. The agreement of $\sigma_{fit}$ with $\sigma$ is good only when the momentum dispersion is smaller than the width of the resonance, estimated by the full width at half maximum (FWHM) $\Delta\tau$ of the Wigner time-delay $\tau$ $\Delta p=0.19$ vs $\Delta\tau=0.24$ in Fig. 3(a)]. Otherwise, $\sigma_{fit}$ is significantly smaller than $\sigma$, suggesting that less entanglement is generated than might be expected. This is readily understood by more momentum components being non-resonant for larger $\Delta p$. For the wave packet with $\Delta p=0.25$, $\sigma_{fit}$ retains a similar shape as $\sigma(p_{0})$, even though the fit parameters $\gamma,\beta$, which were calibrated at $V_{0}=8$, no longer apply here. For even larger $\Delta p$ ($\Delta p=0.37$), $\sigma_{fit}$ deviates from $\sigma$ significantly near the resonance and its peak shifts from the maximum of cross section ($p_{0}=0.64$) towards the maximum of the time delay ($p_{0}\simeq 0.6$), a behavior similar to Fig. 2(c), indicating the influence of interference. Figure 3(b) extends the analysis of the role of the resonance width to multiple $p$-wave shape resonances, cf. Table 1 for the resonance parameters. For $\Delta p=0.19$, the momentum dispersion is smaller than the resonance width for all resonances, cf. Table 2, and $\sigma_{fit}$ and thus $\mathcal{K}(p_{res})$ show the same dependence on $p_{0}$ as the peak of the cross section $\sigma_{max}$. For broader momentum dispersions (pink circles and blue triangles in Fig. 3(b)), $\sigma_{fit}$ drops below $\sigma$, as the narrowness of resonance reduces entanglement generation. For large momentum dispersion ($\Delta p=0.37$), the largest value of $\sigma_{fit}$ (and thus $\mathcal{K}$) does not occur for the strongest resonance (at $p_{res}=0.64$) which is also the narrowest, but rather at $p_{res}=0.82$, a resonance whose FWHM $\Delta\tau\simeq\Delta p$. We thus find that entanglement generation is dominated by the shape resonance, provided the initial momentum dispersion is not too broad since then presence of too many non-resonant momentum components impedes resonance-enhancement of the entanglement generation.
Prospects for probing entanglement generation in collision experiments
Observables that reveal entanglement through measurable quantities are referred to as entanglement witnesses [15]. Unfortunately, the witnesses commonly used for continuous variables in quantum information cannot successfully identify entanglement in scattering states. For example, the PPT criterion requires a decrease in uncertainty when evolving from separable to entangled states [17, 18], but both $\Delta r$ and $\Delta p$ increase upon scattering. The most direct approach would be to probe the single-particle reduced density matrix [32] to extract $\mathcal{K}$. Velocity map imaging (VMI) provides access to the momentum space wavefunction [33]; and combining it with ionization allows for probing elastic collisions [34, 35]. However, VMI yields the amplitudes but not the phases of the momentum space wavefunction. The latter are needed for full quantum state tomography and require an interferometric approach. This can be realized with pump-probe spectroscopy, as recently demonstrated for electronic motion [36, 37]. In order to detect motional entanglement, pump and probe excitation need to couple to free motion of the particle. While such coupling exists (it is at the basis of laser cooling), the momentum changes are likely too small to be measurable in a combination of an interferometric setup with VMI. Instead of reconstructing the full single-particle density matrices, a more practical approach may be to infer the entanglement generated upon a collision by comparing wave packet widths measured in coincidence and single-particle detection, as suggested for electron-ion entanglement in photoionization [38] and atom-photon entanglement in spontaneous emission [39]. To this end, the original proposal [38, 39] needs to be adapted to smaller mass differences. For the experimental implementation, the measurement of elastic collisions [34, 35] would need to be combined by coincidence detection with double-VMI [40, 41, 42]. While clearly challenging, it requires combination of existing experimental technology.
Conclusions
We have analyzed motional entanglement generated upon the collision of two spinless particles in three-dimensional space. While standard scattering theory is based on plane waves, we have found a time-dependent treatment starting from separated wave packets with finite position and momentum uncertainties to be crucial to capture entanglement generation in general and the role of interferences in particular. Quantifying entanglement via the inverse single-particle purity after scattering, we find a linear relationship between entanglement and the collision cross section for sufficiently narrow initial momentum dispersion, in line with earlier findings for pure $s$-wave scattering and very weak collisions [23]. We predict scattering resonances to maximize entanglement generation, unless the momentum dispersion is much broader than the width of the resonance in which case the signature of the resonance is smeared out. While for momentum dispersions smaller than the resonance width, maximum entanglement is generated at the same collision energy at which the cross section peaks, this shifts to the energy which maximizes the Wigner time delay for momentum dispersions comparable with the resonance width. An important next step will be to extend our framework to particles with internal structure, in order to allow for the description of inelastic and reactive collisions. This will open the way to studying the role of entanglement in chemical reaction dynamics.
Acknowledgements.
Acknowledgments
We would like to thank Ed Narevicius for helpful discussions. Financial support by the Alexander von Humboldt Foundation is gratefully acknowledged. We also would like to thank the HPC service and the Department of Physics at Freie Universität Berlin for computing time at the Curta and Sheldon clusters.
Appendix A Methods and Formulation
The single-particle purity ($\mathcal{P}=1/\mathcal{K}$) is obtained as:
$\mathcal{P}(t)$ $=$ $\int d^{3}\vec{r}_{1}d^{3}\vec{r}_{2}d^{3}\vec{r}_{1}^{\,\prime}d^{3}\vec{r}_{2}^{\,\prime}\,\,\tilde{\Psi}^{*}(\vec{r}_{1},\vec{r}_{2}^{\,\prime},t)\tilde{\Psi}^{*}(\vec{r}_{1}^{\,\prime},\vec{r}_{2},t)\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)\tilde{\Psi}(\vec{r}_{1}^{\,\prime},\vec{r}_{2}^{\,\prime},t)$ (3)
$\tilde{\Psi}$ is the scattering wave-packet, which can be derived from the non-scattering wave-packet $\Psi$}. When ${\Psi}(\vec{r}_{1},\vec{r}_{2},t)$ is expanded in the plane-wave basis, its coefficients are the single-particle wavefunctions in momentum space $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ [Eq. (7)]. The plane-wave basis is free to rotate from the separate-particle coordinates, into CM and relative coordinates, using the fractional masses $\mu_{\alpha}$ of each particle ($\mu_{1}+\mu_{2}=1$):
$\Psi(\vec{r}_{1},\vec{r}_{2},t)$ $=$ $\int d^{3}\vec{P}\int d^{3}\vec{p}\,\,\psi_{01}(\mu_{1}\vec{P}+\vec{p},t)\psi_{02}(\mu_{2}\vec{P}-\vec{p},t)e^{i\vec{P}\cdot(\mu_{1}\vec{r}_{1}+\mu_{2}\vec{r}_{2})}e^{i\vec{p}\cdot(\vec{r}_{1}-\vec{r}_{2})}\qquad\text{no sctt}$ (4)
$\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)$ $=$ $\int d^{3}\vec{P}\int d^{3}\vec{p}\,\,\psi_{01}(\mu_{1}\vec{P}+\vec{p},t)\psi_{02}(\mu_{2}\vec{P}-\vec{p},t)e^{i\vec{P}\cdot(\mu_{1}\vec{r}_{1}+\mu_{2}\vec{r}_{1})}f(\vec{p},\vec{r}_{1}-\vec{r}_{2})\quad\text{scatter}$ (5)
The scattering wave-packet $\tilde{\Psi}$ differs from $\Psi$} only in the relative motion $\vec{r}=\vec{r}_{1}-\vec{r}_{2}$. For plane-wave scattering, $e^{i\vec{p}\cdot\vec{r}}\xrightarrow{sctt}f(\vec{p},\vec{r})$, where $f(\vec{p},\vec{r})$ is the eigenfunction of the full scattering Hamiltonian $H_{0}+V(r)$. For $\tilde{\Psi}$ with finite momentum dispersion, the same replacement applied to each $\vec{p}$ component, while the coefficients of the plane-wave basis keep unchanged, assuming scattering is an adiabatic process. For any isotropic scattering potential $V(r)$ decaying no slower than $r^{-2}$, $f(\vec{p},\vec{r})$ can be expanded into partial wave components as:
[…]
with $P_{l}$ the Legendre polynomials, and
$\psi_{l}(p,r)\xrightarrow{r\rightarrow\infty}\frac{1}{2pr}[e^{2i\delta_{l}(p)}e^{i(pr-\frac{l+1}{2}\pi)}+e^{-i(pr-\frac{l+1}{2}\pi)}]$
There exists a cut-off $L$ beyond which scattering phase-shifts $\delta_{l}(p)\rightarrow 0$ and radial wavefunction $\psi_{l}(p,r)\rightarrow j_{l}(pr)$ the first kind spherical Bessel function, when the effective centripetal barrier $\propto\frac{L(L+1)}{r^{2}}$] large enough to block $V(r)$. For example, for finite range potentials in Table 1, we set $L=6$ in the calculation.
The single-particle momentum distribution $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ ($\alpha=1,2$) are Gaussian functions ($\hbar=1$):
$\psi_{0\alpha}(\vec{p}_{\alpha},t)=\left[\frac{\det A}{(4\pi^{3})^{3}}\right]^{\frac{1}{4}}\exp\left[{-\frac{1}{2}(\vec{p}_{\alpha}-\vec{p}_{0\alpha})^{T}A(\vec{p}_{\alpha}-\vec{p}_{0\alpha})}\right]\exp\left[{-\frac{it}{2m}p_{\alpha}^{2}}\right]\exp\left[{i\vec{p}_{\alpha}\cdot\vec{q}_{0\alpha}}\right]$ (7)
We assume the particles to have identical mass $m$ and dispersion $A=\mathrm{diag}(a_{x},a_{y},a_{z})$, with coaxially arranged momenta and locations $\vec{p}_{01}=-\vec{p}_{02}={p}_{0}\hat{z}$, $\vec{q}_{01}=-\vec{q}_{02}=\frac{{q}_{0}}{2}\hat{z}$, so the scattering state $\tilde{\Psi}$ is separable in the CM and relative coordinates. The temporal evolution is parameterized by $\frac{t}{m}$, as shown in Fig. 1(b), where $\mathcal{K}$ shows an inverse linear correlation with the overlap $F=|\langle\Psi|\tilde{\Psi}\rangle|^{2}$, fitted respectively as 6.58F(10,10,20)+6.57-6.58\,F(10,10,20)+6.57, 12.07F(1,1,20)+8.59-12.07\,F(1,1,20)+8.59, and 6.72F(1,1,2)+4.77-6.72\,F(1,1,2)+4.77. Table. 2 gives the fit parameters for post-scattering entanglement $\mathcal{K}(t\to+\infty)$ to cross-sections $\sigma_{fit}$ in Figs. 2 and 3.
Table 2: Parameters of the fit Eq. (2) for the initial conditions in Figs. 2(b,c) and 3.
$\vec{a}$ (20,20,20) (10,10,20) (10,10,20) (10,10,20) (4,4,20)
$q_{0}$ $q_{0}=5$ $q_{0}=9$ $q_{0}=5$ $q_{0}=0$ $q_{0}=5$
$\gamma$ 49.3 23.3 18.1 16.2 9.16
$\beta$ 1.00 0.91 0.74 0.65 1.25
For the numerical calculations, Eq. (3) is integrated using Gauss-Legendre quadrature, with normalization accuracy $\langle\tilde{\Psi}|\tilde{\Psi}\rangle-1<10^{-2}$. The stabilization of $\mathcal{K}$ is confirmed by integrating in each direction of $\vec{r}_{\alpha}$ using different orders of Gauss-Legendre quadrature. The saturated entanglement shown in Fig. 2 and 3 is calculated with time near $t/m\approx 20+q_{0}/(2p_{0})$.
Appendix B Momentum-Space Analysis
Since the wave packet is expanding and oscillating in real space but not in momentum space, on first glance it might seem easier to calculate the entanglement in momentum space [25, 23, 27], using scattering state $\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)$:
$\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)=\int d^{3}\vec{p}\,\,\psi_{01}\left[\mu_{1}(\vec{p}_{1}+\vec{p}_{2})+\vec{p},t\right]\psi_{02}\left[\mu_{2}(\vec{p}_{1}+\vec{p}_{2})-\vec{p},t\right]\,\mathcal{F}(\vec{p},\mu_{2}\vec{p}_{1}-\mu_{1}\vec{p}_{2})\,$, (8)
where $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ ($\alpha=1,2$) are the Gaussian functions in Eq. (7), $\mathcal{F}(\vec{p},\vec{p}_{12})$ is the Fourier transform of $f(\vec{p},\vec{r})$ [Eq. (6)] and $\vec{p}_{12}=\mu_{2}\vec{p}_{1}-\mu_{1}\vec{p}_{2}$ is the relative momentum after scattering. $\mathcal{F}(\vec{p},\vec{p}_{12})=\delta(\vec{p}-\vec{p}_{12})$ only when no scattering occurs $f(\vec{p},\vec{r})=e^{i\vec{p}\cdot\vec{r}}$]. However, otherwise,
$\mathcal{F}(\vec{p},\vec{p}_{12})=\sum_{l=0}^{\infty}\frac{(2l+1)}{8\pi^{2}}P_{l}(\hat{p}\cdot\hat{p}_{12})\left\{\frac{\pi\delta(p-p_{12})}{p_{12}^{2}}(e^{2i\delta_{l}(p)}+1)+\frac{i}{p_{12}p}\textsl{P}\left[\frac{1}{p-p_{12}}+\frac{(-1)^{l+1}}{p+p_{12}}\right](e^{2i\delta_{l}(p)}-1)\,.\right\}$ (9)
The principal value terms $\textit{P}\,[$...] [43] have been incorrectly ignored by a semi-classical model that treats scattering as a “reflection” of relative momentum [26, 27], effectively assuming $|\vec{p}|=|\vec{p}_{12}|$. This assumption preserves the time-dependency as $exp\,{[it{(p_{1}^{2}+p_{2}^{2})}/{2m}]}$ in $\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)$ during the scattering, which ignores the interaction $V$ and results in time-independent $\mathcal{K}$ [25, 23, 26, 27]. Our calculation shows that neglecting $\textit{P}\,[$...] will compromise normalization $\langle\tilde{\Psi}|\tilde{\Psi}\rangle$, violate energy conservation $\langle\tilde{\Psi}|H|\tilde{\Psi}\rangle$, and and underestimate $\mathcal{K}$ up to a factor of 3 under certain initial conditions comparing to its correct asymptotic value $\mathcal{K}(t\to+\infty)$.
To properly integrate over $\vec{p}$ in Eq. (8) including $\textit{P}\,[$...], however, is challenging: The singularity near $p_{12}$ in $\textit{P}\,[$...] is usually handled by the Residue theorem, assuming the remaining integrand is analytic and vanishes at large distance. However, exponential factors such as Gaussian functions $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ diverge over half of the complex plane [e.g., $e^{-(iR)^{2}}\to+\infty$], blocking the construction of a closed contour $C$, making the Residue theorem no longer advantageous. Therefore, evaluating $\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)$ is not easy as initially appears, but rather tricky and difficult. To avoid these issues and fully capture coherence, we evaluate entanglement in real space through $\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)$. Nevertheless, Eq. (9) remains instructive for elucidating why plane-wave initial states fail to generate entanglement, as we explain next.
Appendix C Why plane waves forbid entanglement generation
Assume that the initial state satisfies any two of the following three conditions: (1) $\Delta{P}=0$. (2) $\Delta{p}=0$. (3) The particles are separable prior to collision. Then the pre-scattering state takes the form:
$\psi_{01}(\vec{p}_{1},t)\psi_{02}(\vec{p}_{2},t)=\frac{1}{{N}}e^{-iE_{0}t}\delta(\vec{p}_{1}-\vec{p}_{01})\delta(\vec{p}_{2}-\vec{p}_{02})=\frac{1}{{N}}e^{-iE_{0}t}\delta(\vec{p}_{1}+\vec{p}_{2}-\vec{P}_{0})\delta(\mu_{2}\vec{p}_{1}-\mu_{1}\vec{p}_{2}-\vec{p}_{0})\,$, (10)
where $N=(2\pi)^{3}\delta(\vec{0})$ is the normalization factor, $\vec{p}_{0\alpha}$ is the initial momentum for particle $\alpha$, and $\vec{p}_{0}=\mu_{2}\vec{p}_{01}-\mu_{1}\vec{p}_{02}$ and $\vec{P}_{0}=\vec{p}_{01}+\vec{p}_{02}$ are the relative and CM momentum. $E_{0}=\frac{p_{01}^{2}}{2m_{1}}+\frac{p_{02}^{2}}{2m_{2}}$ is the eigenenergy, the particle masses $m_{1,2}$ and $\mu_{i}=m_{i}/(m_{1}+m_{2})$ are arbitrary.
The scattering wave-packet $\tilde{\Psi}$ and the single-particle purity therefore become:
$\tilde{\Psi}$ $(\vec{p}_{1},\vec{p}_{2},t)=\frac{1}{N}e^{-iE_{0}t}\delta(\vec{p}_{1}+\vec{p}_{2}-\vec{P}_{0})\mathcal{F}(\vec{p}_{0},\mu_{2}\vec{p}_{1}-\mu_{1}\vec{p}_{2})\,$, (11)
$\mathcal{P}$ $=(2\pi)^{12}\int d^{3}\vec{p}_{1}d^{3}\vec{p}_{2}d^{3}\vec{p}_{1}^{\,\prime}d^{3}\vec{p}_{2}^{\,\prime}\,\,\tilde{\Psi}^{*}(\vec{p}_{1},\vec{p}_{2}^{\,\prime},t)\tilde{\Psi}^{*}(\vec{p}_{1}^{\,\prime},\vec{p}_{2},t)\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)\tilde{\Psi}(\vec{p}_{1}^{\,\prime},\vec{p}_{2}^{\,\prime},t)=\frac{(2\pi)^{9}}{N^{3}}\int d^{3}\vec{p}_{12}|\mathcal{F}(\vec{p}_{0},\vec{p}_{12})|^{4}\,.$ (12)
where $\mathcal{F}(\vec{p}_{0},\vec{p}_{12})$ is given in Eq. (9), and can be decompose into non-scattered and a scattered components: $\mathcal{F}(\vec{p}_{0},\vec{p}_{12})=\delta(\vec{p}_{0}-\vec{p}_{12})+\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p}_{12})$, with the latter including finite terms of partial waves characterized by $e^{2i\delta_{l}(p_{0})}-1]$. For $\vec{p}_{0}=\vec{p}_{12}$, $\mathcal{F}(\vec{p}_{0},\vec{p}_{12})=\delta(\vec{p}_{0}-\vec{p}_{12})\gg\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p}_{12})$, otherwise $\mathcal{F}(\vec{p}_{0},\vec{p}_{12})=\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p}_{12})$. Therefore, separating the integral $\int d^{3}\vec{p}_{12}$ into contributions for $\vec{p}_{12}=\vec{p}_{0}$ and $\vec{p}_{12}\neq\vec{p}_{0}$ yields
$\begin{split}\mathcal{P}&=1+\frac{(2\pi)^{9}}{N^{3}}\int_{\vec{p}_{12}\neq\vec{p}_{0}}d^{3}\vec{p}_{12}|\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p}_{12})|^{4}\geq 1\,.\end{split}$ (13)
Since $\mathcal{P}\leq 1$ by definition, Eq. (13) implies $\mathcal{P}=1$, showing that no entanglement is generated ($\mathcal{K}=1$).